OpenStudy (anonymous):

Evaluate the integral: \[ \int_{0}^{1}\; dt \] The integrand is a vector. It is a multivariable calculus question.

OpenStudy (anonymous):

Use the same method you use when taking the intregral of e^-t and 1/1+1. Taking the integral of vectors is the same thing as taking the integral of functions, just keep them in vector form afterard -Take the integral first of the vector function first; the answer will be a vector as well -Then, plug in the given values 0 and 1 so that you have 2 vector functions (one minus another) -Minus the 2 vectors -THAT resulting vector of number/scalar values is your answer

OpenStudy (anonymous):

Integrating e^(-t) we get -e^(-t) and integrating 1/(t+1) gives ln|t+1|, evaluating we get: \[<e^{-1},\frac{1}{2}>-<1,1>=\left<\frac{1}{e}-1, -\frac{1}{2}\right>\]